How many molecules of ammonia are produced from 2.75*10^(-4) grams of hydrogen? Express answer numerically as the number of molecules.


1
Expert's answer
2021-03-11T08:46:43-0500

N2 + 3H2 → 2NH3

M(H2) = 2.016 g/mol

"n = \frac{m}{M}"

n(H2) "= \frac{2.75 \times 10^{-4}}{2.016} = 1.36 \times 10^{-4} \;mol"

According to the reaction equation:

n(NH3) "= \frac{2}{3}n(H2) = 0.90 \times 10^{-4} \;mol"

A mole of ammonia contains 6.02×1023 molecules.

Proportion:

1 mol – "6.02 \times 10^{23}" molecules

"0.90 \times 10^{-4}" mol – x molecules

"x = \frac{6.02 \times 10^{23} \times 0.90 \times 10^{-4}}{1} = 5.418 \times 10^{19}"

Answer: "5.418 \times 10^{19}" molecules